If temperature is constant at 30°C, and pressure decreases from 1.4atm to 0.1atm, which is phase change?


1
Expert's answer
2021-05-24T07:17:54-0400

P1 / T1 = P2 / T2 .


1.4 / 303 = 0.1 / T2


T2 = 303 × 0.1 / 1.4 k = 21.64 k .





If temperature is constant and pressure will decrease than liquid changes into gas .

Solid changes into liquid .